OpenStudy (anonymous):

what is the correct slope and y-intercept of the equation 2x + 3y = 6?

sam (.sam.):

2x + 3y = 6 3y = 6-2x y=-2x/3+2 m=-2/3 , y-int=2

OpenStudy (anonymous):

Oh ok :D thank you!